# GCoT: Chain-of-Thought Prompt Learning for Graphs

### v1

@@ Abstract.
Graph mining, chain-of-thought, prompt learning, pre-training, few-shot learning.
Corresponding authors.
ccs: Information systems Web miningccs: Information systems Data miningccs: Computing methodologies Learning latent representations
@@ 1. Introduction
Recently, Chain-of-Thought (CoT) prompting has demonstrated significant advancement in the field of natural language processing (NLP) (Wei et al., 2022; Wang et al., 2023a; Chu et al., 2023), which mimics the logical process a person may employ to solve a task. Instead of directly providing an answer, CoT prompting decomposes a problem into several steps, guiding the pre-trained language model to follow these steps that lead to the final answer. For example, given the math question in Fig. 1(a), a language model processes with CoT (e.g., initial quantity, eaten quantity, and subtraction) to reach the final answer of 4.
@@
However, the vast potential of CoT prompt learning remains unexplored on pre-trained graph models. Graphs can capture the interactions between entities in a wide range of domains, exhibiting a non-linear topological structure, such as molecular graphs (Wang et al., 2023b; Lee et al., 2023), citation networks (Kanakia et al., 2019; Xiong et al., 2017), and social networks (Ji et al., 2023; Zhang et al., 2023). Conventional approaches typically retrain graph neural networks (GNNs) (Kipf and Welling, 2017; Veličković et al., 2018) or graph transformers (Ying et al., 2021; Yun et al., 2019) for each specific task in an end-to-end supervised manner, relying on abundant labeled data. More recent pre-training methods (Velickovic et al., 2019; You et al., 2020) learn task-invariant, general properties from unlabeled graphs through self-supervised pretext tasks and are then fine-tuned via task-specific labels to adapt to various downstream applications (Hu et al., 2020a; Qiu et al., 2020). To further narrow the gap between pre-training and downstream tasks, prompt learning (Yu et al., 2024a; Liu et al., 2023; Fang et al., 2024) has emerged as a low-resource alternative. They unify the objectives of pre-training and downstream tasks using the same template and employ a lightweight prompt to modify either the input features or hidden embeddings of the pre-trained graph encoder, while keeping the pre-trained weights frozen, as shown in Fig. 1(b). A contemporary work (Jin et al., 2024) leverages graphs to guide reasoning; however, it is built on the CoT mechanism from NLP and relies on textual data to construct thoughts, precluding its application to general text-free graphs. For text-free graphs, existing graph learning methods—including supervised, pre-training, and prompting approaches—produce a “final answer” in a single inference step, which may limit the refinement of the final prediction. In this work, we explore the following question: Would introducing additional inference steps in a CoT style enhance the ability of pre-trained graph models to refine their predictions?
@@
First, what should be the inference steps and thoughts for a graph task? In NLP, a thought is defined as a short instruction that reflects a single reasoning step, with each intermediate textual answer serving as a thought (Chu et al., 2023). However, in general text-free graphs, we cannot directly leverage text as prompts or thoughts. In this work, we aim to improve the inference capability of a pre-trained graph model by incorporating additional steps to refine the answer. We design an inference step with three substages: prompt-based inference, thought construction, and prompt learning. For prompt-based inference, we feed the input graph for the downstream task, along with the prompts, into a pre-trained graph encoder. Then, we construct a “thought” by fusing embeddings from each layer of the pre-trained encoder to capture the current working state with varying levels of graph topological knowledge (Kipf and Welling, 2017). Lastly, the thought is used to learn a set of prompts that guide the next step.
@@
Second, how can we leverage a “thought” to learn prompts and guide the next-step inference? In NLP, chain-of-thought (CoT) prompting is typically implemented by appending specific phrases such as “let’s think step by step” or by providing few-shot CoT examples (Wei et al., 2022; Feng et al., 2024). Then, following a given prompt template, the language model generates new thoughts based on the query and prior thoughts, which in turn facilitate the next reasoning step. In our work, the absence of textual data prevents us from explicitly guiding the next step. Moreover, since each node in a graph exhibits unique characteristics, inference may benefit from node-specific prompts. Thus, we propose a thought-conditioned prompt learning method to guide the next inference step. Specifically, inspired by conditional prompt learning (Zhou et al., 2022), we generate a unique prompt for each node via a conditional network (condition-net), which is conditioned on the node-specific element of the previously constructed thought. The generated prompts then feed into the graph encoder to initiate and guide the next step, repeating the process.
@@
In summary, the contributions of this work are fourfold. (1) We propose GCoT, a Graph CoT prompting approach to guide pre-trained graph models to perform step-by-step inference. To the best of our knowledge, this is the first exploration of CoT prompting on text-free graphs. (2) We design an inference step with three substages: prompt-based inference, thought construction, and thought-conditioned prompt learning. In particular, a thought is constructed by fusing embeddings from each layer of the graph encoder to capture fine-grained topological knowledge in each step. (3) We employ a condition-net to generate node-specific prompts based on the previous thought, enabling the model to perform inference for the downstream task through a step-by-step, node-specific adaptation. (4) We conduct extensive experiments on eight benchmark datasets, demonstrating the superior performance of GCoT compared to a suite of state-of-the-art methods.
@@ 2. Related Work
Chain-of-Thought prompting. Chain-of-Thought (CoT) prompting has emerged as a groundbreaking technique in natural language processing (NLP), empowering language models to address complex reasoning tasks by producing intermediate reasoning steps (Wei et al., 2022; Wang et al., 2023a; Gao et al., 2023). By breaking down problems into a sequence of logical steps, CoT emulates human-like thought processes, leading to significant improvements in model performance on tasks requiring structured reasoning (Yao et al., 2024; Feng et al., 2024). Despite its remarkable success in the NLP domain, the potential of CoT for graphs remains largely unexplored. A contemporary work, GraphCoT (Jin et al., 2024), leverages the inherent relational information in text-attributed graphs (Yan et al., 2023; Wen and Fang, 2023) to guide the reasoning process. However, GraphCoT primarily focuses on question-answering tasks for natural language and cannot be extended to general graphs lacking textual descriptions.
Graph representation learning. GNNs (Kipf and Welling, 2017; Veličković et al., 2018) are dominant technique for graph representations learning. They generally update nodes embeddings iteratively by aggregating information from their local neighborhoods based on a message-passing mechanism (Xu et al., 2019; Hamilton et al., 2017). Despite their success, GNNs often demand substantial amounts of task-specific labeled data and necessitate retraining for each new task, limiting their flexibility and scalability. Recently, researchers have extensively investigated pre-training techniques for graphs (Kipf and Welling, 2016; Hu et al., 2020b, a; Lu et al., 2021). These approaches involve pre-training a graph encoder using self-supervised objectives, and then adapt the pre-trained knowledge to downstream tasks. However, a significant gap exists between the objectives of pre-training and those of downstream tasks, resulting in suboptimal performance.
Graph prompt learning. First proposed in NLP, prompt learning has emerged as a powerful framework for bridging the gap between pre-training and downstream tasks (Brown et al., 2020; Liu et al., 2021b; Lester et al., 2021). Recently, this paradigm has been extended to the graph domain as a compelling alternative to fine-tuning approaches (Liu et al., 2023; Sun et al., 2022; Yu et al., 2024c, d). These methods typically utilize a universal template to align pre-training and downstream tasks, followed by task-specific prompts that facilitate seamless adaptation to downstream tasks while keeping the pre-trained model frozen. However, current graph prompt learning methods directly produce a final answer in a single step, resulting in insufficient refinement to the answer.
@@ 3. Preliminaries
Graph. A graph is defined as $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges. The nodes are associated with a feature matrix $\mathbf{X}\in\mathbb{R}^{|V|\times d}$, where $\mathbf{x}_{v}\in\mathbb{R}^{d}$ is a row of $\mathbf{X}$ representing the feature vector for node $v\in V$. For a collection of multiple graphs, we define it as $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$.
@@
where $\textsc{MP}(\cdot)$ is the message passing function, and $\theta^{l}$ represents the learnable parameters of the graph encoder at layer $l$. The initial embedding matrix, $\mathbf{H}^{0}$, is the input feature matrix, i.e., $\mathbf{H}^{0}=\mathbf{X}$. The output after a total of $L$ layers is then $\mathbf{H}^{L}$; for brevity we simply write $\mathbf{H}$. We abstract the multi-layer encoding process as
@@
where $\{\mathbf{H}^{1},\mathbf{H}^{2},\cdots,\mathbf{H}^{L}\}$ denotes the embedding matrix of the each layer of the graph encoder, respectively. $\Theta=(\theta^{1},\ldots,\theta^{L})$ is the collection of weights across the layers.
Pre-training. As stated by prior studies (Yu et al., 2024d, 2025), all contrastive pre-training task on graphs (Liu et al., 2023; Velickovic et al., 2019; You et al., 2020) can be unified under the task template of similarity calculation. Formally, the unified pre-training objective is defined as follows:
(3) $\mathcal{L}(\Theta)=-\sum_{o\in\mathcal{T}_{\text{pre}}}\ln\frac{\sum_{a\in Pos_{o}}\exp(\text{sim}(\mathbf{h}_{a},\mathbf{h}_{o})/\tau)}{\sum_{b\in Neg_{o}}\exp(\text{sim}(\mathbf{h}_{b},\mathbf{h}_{o})/\tau)},$
where $Pos_{o}$ and $Neg_{o}$ denote the sets of positive and negative samples for a target instance $o$, respectively. $\mathbf{h}_{o}$ represents the embedding of the target instance, while $\mathbf{h}_{a}$ and $\mathbf{h}_{b}$ correspond to the embeddings of positive and negative samples. The hyperparameter $\tau$ controls the temperature scaling in the similarity computation. In our framework, we follow previous work (Liu et al., 2023; Yu et al., 2024d) by employing similarity calculation as the task template and using link prediction as the pre-training task.
Problem definition. In this work, we explore chain-of-thought (CoT) prompt learning framework for text-free graphs. We focus on two widely used tasks in graph learning: node classification and graph classification, in few-shot scenarios. For node classification, given a graph $G=(V,E)$ with a set of node classes $Y$, each node $v_{i}\in V$ is associated with a label $y_{i}\in Y$. In contrast, graph classification considers a collection of graphs $\mathcal{G}$, where each graph $G_{i}\in\mathcal{G}$ is assigned a class label $Y_{i}\in Y$. In the few-shot setting, only $m$ labeled examples per class are available (e.g., $m\leq 10$), a paradigm defined as $m$-shot classification (Liu et al., 2023; Yu et al., 2024d).
@@ 4. Chain-of-Thought Graph Prompt Learning
In this section, we propose our model, GCoT, starting with an overview of its framework. Then we detail its core components and conclude with a complexity analysis of the algorithm.
@@ 4.1. Overall framework
We illustrate the overall framework of GCoT in Fig. 2, consisting of two key stages: pre-training and CoT prompting. First, we pre-train a graph encoder as shown in Fig. 2(a). Details of pre-training are provided in Sect. 3. Second, given a pre-trained graph encoder, to guide the model take additional inference step before finalizing predictions, we propose CoT prompting, as shown in Fig. 2(c). Specifically, we design an inference step with three substages: prompt-based inference, thought construction, and thought-conditioned prompt learning. We first feed the prompt modified query graph into the pre-trained graph encoder. Then we construct a thought by fusing embeddings of all hidden layers of the pre-trained graph encoder. Conditioned on the thought from the previous step, we employ a condition-net to generate a series of node-specific conditional prompts that capture individualized learning patterns for nodes in a fine-grained and parameter-efficient manner, as shown in in Fig. 2(b). We repeat the inference steps until obtain a final answer, as shown in Fig. 2(c). Moreover, to enhance alignment between downstream tasks and the pre-training objective, we incorporate an initial prompt to adjust features or embeddings, same as prior graph prompting methods (Liu et al., 2023; Fang et al., 2024).
@@ 4.2. Chain-of-Thought Prompting
To bridge the gap between the objectives between pre-training and downstream tasks, we adhere to previous graph prompting methods, leveraging an initial prompt to modify node features (Sun et al., 2022; Fang et al., 2024) or output embeddings (Liu et al., 2023; Yu et al., 2025), as shown in Fig. 2(c). Note that our framework is fully compatible with any standard graph prompting techniques and can further enhance their performance, as demonstrated in Sect. 5.5. In our experiments, we employ an attention mechanism to generate prompts that modify the output embedding of the final step, as detailed in the Prompt tuning section.
Prompt-based inference For the $k$-th inference step, we first feed the prompt modified query graph into the pre-trained encoder:
(4) $\{\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}\}=\textsc{GraphEncoder}(\mathbf{X}_{k},G;\Theta_{0})$
where $\Theta_{0}$ is the pre-trained parameters in the graph encoder, $\mathbf{X}_{k}$ is the modified feature by prompts, which will be introduced in the Thought conditioned prompt learning section.
Thought construction To leverage the hierarchical knowledge across multiple layers of the graph encoder, we construct the thought by fusing the hidden embeddings from each layer of the pre-trained graph encoder as follows:
(5) $\mathbf{T}_{k}=\mathtt{Fuse}(\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}),$
where $\mathbf{H}_{k}^{l}$ denotes the hidden embedding from the $l$-th layer during the $k$-th inference step. The $\mathtt{Fuse}(\cdot)$ function can be implemented in various type. In our experiments, we employ weighted summation as the fusion function:

### v2

@@ Abstract.
Graph learning, chain-of-thought, prompt learning.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/25/08doi: 10.1145/3711896.3736974ccs: Information systems Data mining
KDD Availability Link:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15501903.
@@ 1. Introduction
Recently, Chain-of-Thought (CoT) prompting has demonstrated significant advancement in the field of natural language processing (NLP) (Wei et al., 2022; Wang et al., 2023a; Chu et al., 2023), which mimics the logical process a human may employ to solve a task. Instead of directly providing an answer, CoT prompting decomposes a problem into several steps, guiding the pre-trained language model to follow these steps that lead to the final answer. For example, given the math question in Fig. 1(a), a language model utilizes CoT prompting (e.g., initial quantity, eaten quantity, and subtraction) to reach the final answer of 4.
@@
However, the vast potential of CoT prompt learning remains unexplored on pre-trained graph models. Graphs can capture the interactions between entities in a wide range of domains, exhibiting a non-linear topological structure, such as molecular graphs (Wang et al., 2023b; Lee et al., 2023; Li et al., 2024a), citation networks (Kanakia et al., 2019; Xiong et al., 2017; Li et al., 2024b), and social networks (Ji et al., 2023; Zhang et al., 2023). Conventional approaches typically retrain graph neural networks (GNNs) (Kipf and Welling, 2017; Veličković et al., 2018) or graph transformers (Ying et al., 2021; Yun et al., 2019) for each specific task in an end-to-end supervised manner, relying on abundant labeled data. More recent pre-training methods (Velickovic et al., 2019; You et al., 2020) learn task-invariant, general properties from unlabeled graphs through self-supervised pretext tasks and are then fine-tuned via task-specific labels to adapt to various downstream applications (Hu et al., 2020a; Qiu et al., 2020). To further narrow the gap between pre-training and downstream tasks, prompt learning (Yu et al., 2024b; Liu et al., 2023; Fang et al., 2024) has emerged as a low-resource alternative. They unify the objectives of pre-training and downstream tasks using the same template and employ a lightweight prompt to modify either the input features or hidden embeddings of the pre-trained graph encoder, while keeping the pre-trained weights frozen, as shown in Fig. 1(b). A contemporary work (Jin et al., 2024) leverages graphs to guide reasoning; however, it is built on the CoT mechanism from NLP and relies on textual data to construct thoughts, precluding its application to general text-free graphs. For text-free graphs, existing graph learning methods—including supervised, pre-training, and prompting approaches—produce a “final answer” in a single inference step, which may limit the refinement of the final prediction. In this work, we explore the following question: Would introducing additional inference steps in a CoT style enhance the ability of pre-trained graph models to refine their predictions?
@@
First, what should be the inference steps and thoughts for a graph task? In NLP, a thought is defined as a short instruction that reflects a single reasoning step, with each intermediate textual answer serving as a thought (Chu et al., 2023). However, in general text-free graphs, we cannot directly leverage text as prompts or thoughts. In this work, we aim to improve the inference capability of a pre-trained graph model by incorporating additional steps to refine the answer. We design an inference step with three substages: prompt-based inference, thought construction, and prompt learning. For prompt-based inference, we feed the input graph for the downstream task, along with some prompts, into a pre-trained graph encoder. Then, we construct a “thought” by fusing embeddings from each layer of the pre-trained encoder to capture the current working state with varying levels of topological knowledge (Kipf and Welling, 2017). Lastly, the thought is used to learn a set of prompts that guide the next step.
@@
Second, how can we leverage a “thought” to learn prompts and guide the next-step inference? In NLP, CoT prompting is typically implemented by appending specific phrases such as “let’s think step by step” or by providing few-shot CoT examples (Wei et al., 2022; Feng et al., 2024). Then, following a given prompt template, the language model generates new thoughts based on the query and prior thoughts, which in turn facilitate the next reasoning step. In our work, the absence of textual data prevents us from explicitly guiding the next step. Moreover, since each node in a graph exhibits unique characteristics, inference may benefit from node-specific prompts. Thus, we propose a thought-conditioned prompt learning method to guide the next inference step. Specifically, inspired by conditional prompt learning (Zhou et al., 2022), we generate a unique prompt for each node via a conditional network (condition-net), which is conditioned on the node-specific element of the previously constructed thought. The generated prompts then feed into the graph encoder to initiate and guide the next step, repeating the process.
@@
In summary, the contributions of this work are fourfold. (1) We propose GCoT, a Graph CoT prompting approach to guide pre-trained graph models to perform step-by-step inference. To the best of our knowledge, this is the first exploration of CoT-style prompting on text-free graphs. (2) We design an inference step with three substages: prompt-based inference, thought construction, and thought-conditioned prompt learning. In particular, a thought is constructed by fusing embeddings from each layer of the graph encoder to capture fine-grained topological knowledge in each step. (3) We employ a condition-net to generate node-specific prompts based on the previous thought, enabling the model to perform inference for the downstream task through a step-by-step, node-specific adaptation. (4) We conduct extensive experiments on eight benchmark datasets, demonstrating the superior performance of GCoT compared to a suite of state-of-the-art methods.
@@ 2. Related Work
In this section, we briefly review related work on CoT prompting, graph learning, and graph prompt learning.
Chain-of-Thought prompting. Chain-of-Thought (CoT) prompting has emerged as a groundbreaking technique in NLP, empowering language models to address complex reasoning tasks by producing intermediate reasoning steps (Wei et al., 2022; Wang et al., 2023a; Gao et al., 2023). By breaking down problems into a sequence of logical steps, CoT emulates human-like thought processes, leading to significant improvements in model performance on tasks requiring structured reasoning (Yao et al., 2024; Feng et al., 2024). Despite its remarkable success in NLP applications, the potential of CoT prompting for graphs remains largely unexplored. A contemporary work, GraphCoT (Jin et al., 2024), leverages the inherent relational information in text-attributed graphs (Yan et al., 2023; Wen and Fang, 2023) to guide the reasoning process. However, GraphCoT primarily focuses on question-answering tasks for natural language and cannot be extended to general text-free graphs that lack textual descriptions.
Graph representation learning. GNNs (Kipf and Welling, 2017; Veličković et al., 2018; Yu et al., 2023; Jiang et al., 2024, 2023b) are the dominant technique for graph representations learning. They generally update node embeddings iteratively by aggregating information from their local neighborhoods based on a message-passing mechanism (Xu et al., 2019; Hamilton et al., 2017; Li et al., 2024c). Despite their success, GNNs often demand substantial amounts of task-specific labeled data and necessitate retraining for each new task, limiting their flexibility and scalability. Recently, researchers have extensively investigated pre-training techniques for graphs (Kipf and Welling, 2016; Hu et al., 2020b, a; Lu et al., 2021; Liu et al., 2025; Jiang et al., 2023a). These approaches involve pre-training a graph encoder using self-supervised objectives, and then adapt the pre-trained knowledge to downstream tasks. However, a significant gap exists between the objectives of pre-training and those of downstream tasks, resulting in suboptimal performance.
Graph prompt learning. First proposed in NLP, prompt learning has emerged as a powerful framework for bridging the gap between pre-training and downstream tasks (Brown et al., 2020; Liu et al., 2021b; Lester et al., 2021). Recently, this paradigm has been extended to the graph domain as a compelling alternative to fine-tuning approaches (Liu et al., 2023; Sun et al., 2022; Yu et al., 2024a, c, d). These methods typically utilize a universal template to align pre-training and downstream tasks, followed by task-specific prompts that facilitate seamless adaptation to downstream tasks while keeping the pre-trained model frozen. However, current graph prompt learning methods directly produce a final answer in a single step, resulting in insufficient refinement to the answer (Yu et al., 2024e, 2025a; Fang et al., 2025).
@@ 3. Preliminaries
In this section, we present the background and preliminaries relevant to our work.
Graph. A graph is defined as $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges. The nodes are associated with a feature matrix $\mathbf{X}\in\mathbb{R}^{|V|\times d}$, where $\mathbf{x}_{v}\in\mathbb{R}^{d}$ is a row of $\mathbf{X}$ representing the feature vector for node $v\in V$. For a collection of multiple graphs, we denote it as $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$.
@@
where $\textsc{MP}(\cdot)$ is the message passing function, and $\theta^{l}$ represents the learnable parameters of the graph encoder at layer $l$. The initial embedding matrix, $\mathbf{H}^{0}$, is the input feature matrix, i.e., $\mathbf{H}^{0}=\mathbf{X}$. The output after a total of $L$ layers is then $\mathbf{H}^{L}$; for brevity we may simply write $\mathbf{H}$. We abstract the multi-layer encoding process as
@@
where $\{\mathbf{H}^{1},\mathbf{H}^{2},\cdots,\mathbf{H}^{L}\}$ denotes the embedding matrix of each layer of the graph encoder, respectively. $\Theta=(\theta^{1},\ldots,\theta^{L})$ is the collection of weights across the layers.
Pre-training. As prior studies (Yu et al., 2024c, 2025c) have suggested, mainstream contrastive pre-training tasks on graphs (Liu et al., 2023; Velickovic et al., 2019; You et al., 2020) can be unified under the task template of similarity calculation. Formally, the unified pre-training objective is defined as follows:
(3) $\mathcal{L}(\Theta)=-\sum_{o\in\mathcal{T}_{\text{pre}}}\ln\frac{\sum_{a\in\text{Pos}_{o}}\exp(\text{sim}(\mathbf{h}_{a},\mathbf{h}_{o})/\tau)}{\sum_{b\in\text{Neg}_{o}}\exp(\text{sim}(\mathbf{h}_{b},\mathbf{h}_{o})/\tau)},$
where $\text{Pos}_{o}$ and $\text{Neg}_{o}$ denote the sets of positive and negative samples for a target instance $o$, respectively. $\mathbf{h}_{o}$ represents the embedding of the target instance, while $\mathbf{h}_{a}$ and $\mathbf{h}_{b}$ correspond to the embeddings of positive and negative samples. The hyperparameter $\tau$ controls the temperature scaling of the similarity function $\text{sim}(\cdot,\cdot)$. In our framework, we follow previous work (Liu et al., 2023; Yu et al., 2024c) by employing similarity calculation as the task template and using link prediction as the pre-training task.
Problem definition. In this work, we explore a CoT prompt learning framework for text-free graphs. We focus on two widely used tasks in graph learning: node classification and graph classification, in few-shot scenarios. For node classification, given a graph $G=(V,E)$ with a set of node classes $Y$, each node $v_{i}\in V$ is associated with a label $y_{i}\in Y$. In contrast, graph classification considers a collection of graphs $\mathcal{G}$, where each graph $G_{i}\in\mathcal{G}$ is assigned a class label $Y_{i}\in Y$. In the few-shot setting, only $m$ labeled examples per class are available (e.g., $m\leq 10$), which is referred to as $m$-shot classification (Liu et al., 2023; Yu et al., 2024c).
@@ 4. Chain-of-Thought Graph Prompt Learning
In this section, we present our approach, GCoT, starting with an overview of its framework. Then we detail its core components and conclude with a complexity analysis of the algorithm.
@@ 4.1. Overall Framework
We illustrate the overall framework of GCoT in Fig. 2, which consists of two phases: pre-training and CoT prompting.
First, we pre-train a graph encoder as shown in Fig. 2(a). Details of the pre-training phase are provided in Sect. 3.
Second, given a pre-trained graph encoder, to guide the model with additional inference steps before finalizing its predictions, we propose CoT prompting, as illustrated in Fig. 2(c). Specifically, we design an inference step with three substages: prompt-based inference, thought construction, and thought-conditioned prompt learning. Specifically, in each inference step, we first feed the prompt and query graph into the pre-trained graph encoder. Then, we construct a “thought” by fusing embeddings of all hidden layers of the pre-trained graph encoder. Lastly, as detailed in Fig. 2(b), conditioned on the constructed thought, we employ a condition-net to generate a series of node-specific prompts. These conditional prompts capture individualized patterns for nodes in a fine-grained and parameter-efficient manner. We repeat the inference step one or more times before making the final prediction for the downstream task. It is worth noting that, to better align the downstream task with the pre-training task, we still incorporate a standard prompt similar to those used in prior graph prompting methods (Liu et al., 2023; Fang et al., 2024).
@@ 4.2. Chain-of-Thought Prompting
Next, we introduce the details of each inference step and the standard prompt learning employed in our framework.
4.2.1. Inference step
Each inference step consists of three substages, as described in the following.
Prompt-based inference. In the $k$-th inference step, we first feed the query graph $G=(V,E)$, along with its prompt-modified feature matrix, into the pre-trained encoder:
(4) $\{\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}\}=\textsc{GraphEncoder}(\mathbf{X}_{k},G;\Theta_{0}),$
where $\Theta_{0}$ denotes the frozen pre-trained weights in the graph encoder, and $\mathbf{X}_{k}$ is the node feature matrix modified by the prompts generated in the previous (i.e., $k-1$-th) step. The generation of the prompts and the modification of the feature matrix are elaborated later in thought-conditioned prompt learning. Note that for the first step ($k=1$), the original feature matrix is used without modification, i.e., $\mathbf{X}_{1}=\mathbf{X}$.
Thought construction. To leverage the hierarchical knowledge across multiple layers of the graph encoder, we construct a “thought” by fusing the hidden embeddings from each layer of the pre-trained graph encoder as follows:
(5) $\mathbf{T}_{k}=\textsc{Fuse}(\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}),$
where each $\mathbf{H}_{k}^{l}$ denotes the hidden embedding from the $l$-th layer during the $k$-th inference step. The $\textsc{Fuse}(\cdot)$ function can be implemented in various ways. For example, if all hidden layers of the graph encoder share the same dimensionality, we can adopt a simple weighted summation [^1] This is the case in our experiments. If the dimensionalities differ, linear layers can be potentially used to project them to a common dimension., as follows.